OpenStudy (anonymous):

Add the rational expression. The answer should be in lowest terms. x div x ^{2}+3x+2 + 2x-3div x ^{2} -1

OpenStudy (anonymous):

\[\frac{x^2+3x+2}{x^2-1}=\frac{(x+2)(x+1)}{(x+1)(x-1)}\]x+1 cancle\[\frac{x+2}{x-1}\]

jhonyy9 (jhonyy9):

x 2x-3 x 2x-3 x(x-1) + (x+2)(2x-3) ---------- + ------- = --------- + --------- = ------------------ x*2 +3x +2 x*2 -1 (x+1)(x+2) (x-1)(x+1) (x-1)(x+1)(x+2)